# BEASST: Behavioral Entropic Gradient based Adaptive Source Seeking for Mobile Robots

### v1

@@ Abstract
This paper presents BEASST (Behavioral Entropic Gradient-based Adaptive Source Seeking for Mobile Robots), a novel framework for robotic source seeking in complex, unknown environments. Our approach enables mobile robots to efficiently balance exploration and exploitation by modeling normalized signal strength as a surrogate probability of source location. Building on Behavioral Entropy(BE) with Prelec’s probability weighting function, we define an objective function that adapts robot behavior from risk-averse to risk-seeking based on signal reliability and mission urgency. The framework provides theoretical convergence guarantees under unimodal signal assumptions and practical stability under bounded disturbances. Experimental validation across DARPA SubT and multi-room scenarios demonstrates that BEASST consistently outperforms state-of-the-art methods, achieving 15% reduction in path length and 20% faster source localization through intelligent uncertainty-driven navigation that dynamically transitions between aggressive pursuit and cautious exploration.
@@ I Introduction
This paper addresses active source seeking (SS) in unknown environments via a mobile robot. SS aims to move the robot to a target location based on perceived information from a signal of interest (SoI), such as radio signals, radiation sources, or plumes from hazardous agents [1, 2, 3, 4]. These signals are often affected by environmental obstacles and distance, making target localization challenging across applications including environmental monitoring, search and rescue operations, and hazardous material detection.
Traditionally, SS methods often rely on extremum seeking controllers [5, 6] or employ simplified probabilistic models that assume static uncertainty where every measurement is treated equally trustworthy [1, 7]. Extremum seeking approaches, while providing real-time capabilities, suffer from vanishing gradients in far-field regions and overshooting behavior near sources, limiting their effectiveness in complex signal landscapes. Alternative approaches include bio-inspired methods such as infotaxis [8] for odor source localization, which exploit intermittent signal detection patterns but often lack theoretical convergence guarantees. Multi-robot cooperative strategies [9, 10] have demonstrated promise in distributed scenarios but introduce coordination complexity and communication overhead that can limit scalability. Similarly, discrete planning approaches evaluate information gain at the boundary of known and unknown areas (frontiers), greedily selecting them until sources are found [11, 12]. While entropy-based frameworks like Infosploration [12] use Shannon entropy (SE) to model environmental uncertainty for strategic trajectory planning, they often struggle with global optimality and efficient information gathering in complex environments with varying signal characteristics. The fundamental limitation lies in their uniform treatment of uncertainty across different signal strengths, failing to capture the nuanced decision-making required for effective source seeking. A critical limitation of these conventional methods is their inability to dynamically adapt motion strategies based on perceived signal strength, leading to vanishing gradients or overshooting. Furthermore, diverse SS applications demand flexible policies that can respond to varying environmental conditions and signal characteristics, often requiring extensive manual parameter tuning to balance exploration and exploitation effectively.
@@
  • Behavioral Entropic Gradients (BUGs): A novel SS control framework via localized differential BE, formulated from 2-Wasserstein gradient flow to yield adaptive source-seeking velocity fields with theoretical convergence guarantees.
@@
We validate our framework through several ROS-Unity simulation environments with different configurations and several comparisons with frontier-based SS methods like Infosploration [12], Behaviorsploration  [^1] Is an extension of Infosploration by incorporating BE to evaluate information gain of frontiers during the exploration phase and random frontiers.
@@ II Problem Setting and Preliminaries
Central to BE is Prelec’s probability weighting function [15], $w:[0,1]\rightarrow[0,1]$, defined as:
$w(p)=\textup{e}^{-\beta(-\log(p))^{\alpha}},$ (2)
where $p$ is an ‘objective’ uncertainty measure. [^2] For the differential form of BE, where $p(\bm{\mathbf{x}})$ is a probability density function, special consideration is often required as $p(\bm{\mathbf{x}})$ can take values outside $[0,1]$ [16]. In our application, however, $p(\bm{\mathbf{x}})$ represents a normalized surrogate probability derived from signal strength, which by definition lies within the $[0,1]$ range, ensuring compatibility with $w(p)$. When $\alpha=1$ and $\beta=1$, $\mathcal{H}_{B}$ equates to Shannon entropy. The parameter $\alpha>0$ controls the convexity or concavity of $w$, significantly altering perceived probability: a small $\alpha\approx 0$ makes $w$ concave and over-weights probabilities ($w(p)\gg p$), while $\alpha\to\infty$ under-weights them ($w(p)\ll p$). The parameter $\beta>0$ scales the overall weighting [13]. Thus, Prelec’s function $w$ encodes the robot’s ‘subjective perception’ of uncertainty, making $w(p)$ the perceived uncertainty measure. BE’s parameters ($\alpha,\beta$) can dynamically change the robot’s behavior by altering its perceived notion of uncertainty. For instance, a small $\alpha$ can magnify perceived uncertainty in regions of low objective probability, making the robot more uncertainty-averse in those areas.
@@ III Behavioral Source Seeking
$p(\bm{\mathbf{x}})=\frac{P_{dBm}(\bm{\mathbf{x}})}{P_{dBm}^{\text{max}}},$ (3)
where $P_{dBm}(\bm{\mathbf{x}})$ is the detected signal strength at $\bm{\mathbf{x}}$, and $P_{dBm}^{\text{max}}$ is the maximum and a priori known signal strength that is usually expected at the source location. In our SS, this normalized signal strength function serves as a surrogate probability density function. Its value is almost $1$ at the source location and decreases with distance, reflecting the diminishing signal strength. Our rigorous theoretical development presented below relies on the following assumptions regarding the characteristics of the signal strength and signal reception. It is worth noting that empirical studies carried out in Section VI indicate the method’s robust performance even in more complex real-world scenarios where obstacles and obstructions may lead to localized deviations from these ideal continuity conditions.
@@ IV Behavioral motion policy design for source seeking
$\delta\mathcal{H}_{B}(w,\delta w)$ $=\frac{\text{d}}{\text{d}\epsilon}\mathcal{H}_{B}(w(\mathbf{x})+\delta w(\mathbf{x}))\bigg{|}_{\epsilon=0}$ (7)
@@ IV-A Performance Analysis
where $\bm{\mathbf{\omega}}(t)$ is a bounded disturbance satisfying $\|\bm{\mathbf{\omega}}(t)\|\leq\bar{\omega}$ for some $\bar{\omega}\geq 0$. Let $p(\bm{\mathbf{x}})$ be the normalized measured signal strength defined according to (3), and let Assumption 2 hold. Then, for any $\alpha,\beta>0$, the robot’s position $\bm{\mathbf{x}}_{r}(t)$ will ultimately converge to and remain within a compact ball of radius $R(\bar{\omega})$ centered at the source location $\bm{\mathbf{x}}_{s}$, where $R(\bar{\omega})\to 0$ as $\bar{\omega}\to 0$.
@@
Consider the Lyapunov function $V(\bm{\mathbf{x}}_{r})=-\log(w(p(\bm{\mathbf{x}}_{r})))$ from Theorem IV.1. The time derivative along trajectories of (12) is:
@@
Under Assumption 2, $p(\bm{\mathbf{x}})$ has a unique global maximum at $\bm{\mathbf{x}}_{s}$, which implies $\nabla p(\bm{\mathbf{x}})\neq 0$ for all $\bm{\mathbf{x}}\neq\bm{\mathbf{x}}_{s}$. Since $w(p)$ is strictly increasing with $w^{\prime}(p)>0$ for $p\in(0,1)$, we have $\bm{\mathbf{g}}(\bm{\mathbf{x}}_{r})=\frac{w^{\prime}(p(\bm{\mathbf{x}}_{r}))}{w(p(\bm{\mathbf{x}}_{r}))}\nabla p(\bm{\mathbf{x}}_{r})\neq 0$ for all $\bm{\mathbf{x}}_{r}\neq\bm{\mathbf{x}}_{s}$. By continuity of $\bm{\mathbf{g}}(\bm{\mathbf{x}}_{r})$ and compactness arguments, for any $\delta>0$, there exists $\gamma(\delta)>0$ such that $\|\bm{\mathbf{g}}(\bm{\mathbf{x}}_{r})\|^{2}\geq\gamma(\delta)$, whenever $\|\bm{\mathbf{x}}_{r}-\bm{\mathbf{x}}_{s}\|\geq\delta$. Moreover, $\gamma(\delta)$ is increasing in $\delta$ and $\gamma(\delta)\to 0^{+}$ as $\delta\to 0^{+}$. Substituting into (14), we obtain:
@@
This can be written in the standard ISpS form:
@@
Figure 4: Comparison of convergence steps for BEASST with adaptive tuning versus fixed $\alpha=\beta=1$, which corresponds to use of SE from two distinct starting positions( Scenario 1 and 2).
@@ IV-B Adaptive Parametric Tuner
The crucial reshaping capability and tuning power are lost with Shannon differential entropy ($\alpha=\beta=1$ in BE $\mathcal{H}_{B}$). In this case, the control law (10) simplifies to $\dot{\bm{\mathbf{x}}}_{r}(t)=\frac{1}{p(\bm{\mathbf{x}}_{r}(t))}\nabla p(\bm{\mathbf{x}}_{r}(t))$. This non-adaptive approach leads to significantly slower convergence, especially from areas with low signal strength, unlike our method which enables effective SS even with weaker signals. This property of the tuning mechanism is beneficial for SS in unknown environments, as it expedites the process by allowing a lower signal detection threshold for transitioning from exploration to source seeking. This effect is clearly illustrated in two scenarios in Fig. 4 where the underlying probability field is modeled by a log-normal distribution and the convergence to the source is faster using our adaptive algorithm. We choose log normal in this example to demonstrate the effectiveness of our algorithm even in asymmetric signals that closely match real SoIs, such as acoustic signals.
1 Input: Occupancy map $\mathcal{M}_{D}$, measured signal strength $P_{dBm}(\bm{\mathbf{x}})$, its gradient $\nabla P_{dBm}(\bm{\mathbf{x}})$, maximum expected signal strength $P_{dBm}^{\text{max}}$, number of sources $N_{s}$;
@@ 2 Output: Robot waypoints $\{\bm{\mathbf{x}}_{t+1}\}$;
3 Initialize: $S_{\text{detected}}=0$, $\bm{\mathbf{x}}_{0}$, $\alpha,\beta=1$$\tau$,$t=0$;
4while $S_{\text{detected}}<N_{s}$ do
@@ 3 Initialize: $S_{\text{detected}}=0$, $\bm{\mathbf{x}}_{0}$, $\alpha,\beta=1$$\tau$,$t=0$;
  $p_{t}(\bm{\mathbf{x}})\leftarrow\frac{P_{dBm}(\bm{\mathbf{x}}_{t})}{P_{dBm}^{\text{max}}}$; // Normalize using Eq. 3
@@ 11    $\bm{\mathbf{x}}_{t+1}\leftarrow\bm{\mathbf{x}}_{t}+\gamma\nabla\log(w(p_{t}(\bm{\mathbf{x}})))$;
13        $S_{\text{detected}}\leftarrow S_{\text{detected}}+1$;
14        UpdateSignalMap();
15     end if
@@ 19 $\bm{\mathbf{x}}_{t+1}\leftarrow\text{GPBasedExp}(\mathcal{M}_{D},\text{p(occ(x))})$;
  Update($\mathcal{M}_{D}$) // SLAM Update [23]
20
@@ V Behavioral Source Seeking with Exploration
The robot initially explores the unknown environment. Upon detecting a signal, above a predefined normalized threshold $\tau>0$, it transitions to SS, driven by the gradients (11) from normalized signal strength $p_{t}(\bm{\mathbf{x}})$. This operation is performed by SourceSensing(.) function in Algorithm 1. The gradient is computed using Gradient(.), which naturally guides the robot towards the strongest perceived signal using adaptive $\alpha_{t}$. For practical implementation, a discrete-time update for the robot’s position is given in Line 10 of Algorithm 1, where $\bm{\mathbf{x}}(t)$ is the desired waypoint for the robot to track. When multiple SoIs are detected, the framework strategically marks the locations of other significant (but not immediately pursued) SoIs as ‘virtual frontiers’. Once a current source has been adequately investigated, then the signal map is updated using the UpdateSignalMap function. These virtual frontiers serve as waypoints for future exploration, ensuring comprehensive coverage. If no significant signal is detected, the robot prioritizes exploration, moving towards regions of high mapping uncertainty based on GP-based exploration and Omnimapper [23] updates the map (Update Function). This adaptive switching allows the robot to efficiently locate and map multiple sources.

### v2

@@ Abstract
This paper presents BEASST (Behavioral Entropic Gradient-based Adaptive Source Seeking for Mobile Robots), a novel framework for robotic source seeking in complex, unknown environments. Our approach enables mobile robots to efficiently balance exploration and exploitation by modeling normalized signal strength as a surrogate probability of source location. Building on Behavioral Entropy (BE) with Prelec’s probability weighting function, we define an objective function that adapts robot behavior from risk-averse to risk-seeking based on signal reliability and mission urgency. The framework provides theoretical convergence guarantees under unimodal signal assumptions and practical stability under bounded disturbances. Experimental validation across DARPA SubT and multi-room scenarios demonstrates that BEASST consistently outperforms state-of-the-art methods and exhibits strong robustness to noisy gradient estimates while maintaining convergence. BEASST achieved 15% reduction in path length and 20% faster source localization through intelligent uncertainty-driven navigation that dynamically transitions between aggressive pursuit and cautious exploration.
@@ I Introduction
This paper addresses active source seeking (SS) in unknown environments via a mobile robot. SS aims to move the robot to a target location based on perceived information from a signal of interest (SoI), such as radio signals, radiation sources, or plumes from hazardous agents [1, 2, 3, 4]. These signals are often affected by environmental obstacles and distance, making target localization challenging across applications including environmental monitoring, search and rescue operations, and hazardous material detection. Traditionally, SS methods often rely on extremum seeking controllers [5], flow field tracking [6], or employ simplified probabilistic models that assume static uncertainty where every measurement is treated equally trustworthy [1, 7]. Extremum-seeking approaches, while providing real-time capabilities, suffer from vanishing gradients in far-field regions and overshooting behavior near sources, limiting their effectiveness in complex signal landscapes. Alternative approaches include bio-inspired methods such as infotaxis [8] for odor source localization, which exploit intermittent signal detection patterns but often lack theoretical convergence guarantees. Multi-robot cooperative strategies [9, 10, 11] have demonstrated promise in distributed scenarios but introduce coordination complexity and communication overhead that can limit scalability. Similarly, discrete planning approaches evaluate information gain at the boundary of known and unknown areas (frontiers), greedily selecting them until sources are found [12, 13]. While entropy-based frameworks like Infosploration [13] use Shannon entropy (SE) to model environmental uncertainty for strategic trajectory planning, they often struggle with global optimality and efficient information gathering in complex environments with varying signal characteristics. The fundamental limitation lies in their uniform treatment of uncertainty across different signal strengths, failing to capture the nuanced decision-making required for effective source seeking. A critical limitation of these conventional methods is their inability to dynamically adapt motion strategies based on perceived signal strength, leading to vanishing gradients or overshooting. Furthermore, diverse SS applications demand flexible policies that can respond to varying environmental conditions and signal characteristics, often requiring extensive manual parameter tuning to balance exploration and exploitation effectively.
@@
  • Behavioral Entropic Gradients (BUGs): A novel SS control framework via localized differential BE, formulated from 2-Wasserstein gradient flow to yield adaptive source-seeking velocity fields with theoretical convergence and stability guarantees under bounded Gaussian noise.
@@
We validate our framework through several ROS-Unity simulation environments with different configurations and several comparisons with frontier-based SS methods like Infosploration [13], Behaviorsploration  [^1] Is an extension of Infosploration by incorporating BE to evaluate information gain of frontiers during the exploration phase. and random frontiers. Moreover, we show BEASST exhibits strong robustness to noisy gradients in these environments, and maintains convergence and performance under bounded Gaussian disturbances consistent with our theoretical stability guarantees.
@@ II Problem Setting and Preliminaries
Central to BE is Prelec’s probability weighting function [16], $w:[0,1]\rightarrow[0,1]$, defined as: $w(p)=\textup{e}^{-\beta(-\log(p))^{\alpha}},$ where $p$ is an ‘objective’ uncertainty measure. [^2] For the differential form of BE, where $p(\bm{\mathbf{x}})$ is a probability density function, special consideration is often required as $p(\bm{\mathbf{x}})$ can take values outside $[0,1]$ [17]. In our application, however, $p(\bm{\mathbf{x}})$ represents a normalized surrogate probability derived from signal strength, which by definition lies within the $[0,1]$ range, ensuring compatibility with $w(p)$. When $\alpha=1$ and $\beta=1$, $\mathcal{H}_{B}$ equates to Shannon entropy. The parameter $\alpha>0$ controls the convexity or concavity of $w$, significantly altering perceived probability: a small $\alpha\approx 0$ makes $w$ concave and over-weights probabilities ($w(p)\gg p$), while $\alpha\to\infty$ under-weights them ($w(p)\ll p$). The parameter $\beta>0$ scales the overall weighting [14]. Thus, Prelec’s function $w$ encodes the robot’s ‘subjective perception’ of uncertainty, making $w(p)$ the perceived uncertainty measure. BE’s parameters ($\alpha,\beta$) can dynamically change the robot’s behavior by altering its perceived notion of uncertainty. For instance, a small $\alpha$ can magnify perceived uncertainty in regions of low objective probability, making the robot more uncertainty-averse in those areas.
@@ III Behavioral Source Seeking
$p(\bm{\mathbf{x}})=\phi({P_{dBm}(\bm{\mathbf{x}})}),\quad\phi:\mathbb{R}\rightarrow(0,1]$ (2)
where $P_{dBm}(\bm{\mathbf{x}})$ is the detected signal strength at $\bm{\mathbf{x}}$, where $\phi(\cdot)$ is a continuously differentiable and strictly increasing normalizing function. This general formulation encompasses various normalizations, where one could choose $(\phi(s)=\frac{s}{Z})$, $(\phi(s)=s/P_{dBm}^{\max})$ or any increasing function. In our SS, this normalized signal strength function serves as a surrogate probability density function. Its value is almost $1$ at the source location and decreases with distance, reflecting the diminishing signal strength. Our rigorous theoretical development presented below relies on the following assumptions regarding the characteristics of the signal strength and signal reception. It is worth noting that empirical studies carried out in Section VI indicate the method’s robust performance in complex real-world scenarios where obstacles and noise may lead to localized deviations from ideal continuity conditions.
@@ IV Behavioral motion policy design for source seeking
$\delta\mathcal{H}_{B}(w,\delta w)$ $=\frac{\text{d}}{\text{d}\epsilon}\mathcal{H}_{B}(w(\mathbf{x})+\delta w(\mathbf{x}))\bigg|_{\epsilon=0}$
@@ IV-A Performance Analysis
where $\bm{\mathbf{\omega}}(t)$ is a bounded disturbance satisfying $\|\bm{\mathbf{\omega}}(t)\|\leq\bar{\omega}$ for some $\bar{\omega}\geq 0$. Let $p(\bm{\mathbf{x}})$ be the normalized signal strength defined in  (2), and let Assumption 2 (unimodal signal with unique maximum) hold. Then, for any $\alpha,\beta>0$, the robot’s position $\bm{\mathbf{x}}_{r}(t)$ will ultimately converge to and remain within a compact ball of $\mathcal{B}_{R(\bar{\omega})}(\bm{\mathbf{x}}_{s})$ of radius $R(\bar{\omega})$ centered at the source location $\bm{\mathbf{x}}_{s}$, where $R(\bar{\omega})\to 0$ as $\bar{\omega}\to 0$.
@@
Consider the Lyapunov function $V(\bm{\mathbf{x}}_{r})=-\log(w(p(\bm{\mathbf{x}}_{r})))$ as in the noisy-free case from Theorem IV.1. Differentiating $V$ along trajectories  (10) gives:
@@
Under Assumption 2, $p(\bm{\mathbf{x}})$ has a unique global maximum at $\bm{\mathbf{x}}_{s}$, which implies $\nabla p(\bm{\mathbf{x}})\neq 0$ for all $\bm{\mathbf{x}}\neq\bm{\mathbf{x}}_{s}$. Since $w(p)$ is strictly increasing with $w^{\prime}(p)>0$ for $p\in(0,1)$, the composite gradient $\bm{\mathbf{g}}(\bm{\mathbf{x}}_{r})=\frac{w^{\prime}(p(\bm{\mathbf{x}}_{r}))}{w(p(\bm{\mathbf{x}}_{r}))}\nabla p(\bm{\mathbf{x}}_{r})\neq 0$ for all $\bm{\mathbf{x}}_{r}\neq\bm{\mathbf{x}}_{s}$. By continuity of $\bm{\mathbf{g}}(\bm{\mathbf{x}}_{r})$ and compactness arguments, for any $\delta>0$, there exists $\gamma(\delta)>0$ such that $\|\bm{\mathbf{g}}(\bm{\mathbf{x}}_{r})\|^{2}\geq\gamma(\delta)$, whenever $\|\bm{\mathbf{x}}_{r}-\bm{\mathbf{x}}_{s}\|\geq\delta$. Moreover, $\gamma(\delta)$ is increasing in $\delta$ and $\gamma(\delta)\to 0^{+}$ as $\delta\to 0^{+}$. Substituting into (12), we obtain:
@@
This has the standard ISpS structure:
@@
Figure 4: Comparison of convergence steps for BEASST with adaptive tuning, and its noisy variant versus fixed $\alpha=\beta=1$, which corresponds to use of SE from two distinct starting positions( Scenario 1 and 2).
@@ IV-B Adaptive Parametric Tuner
The crucial reshaping capability and tuning power are lost with Shannon differential entropy ($\alpha=\beta=1$ in BE $\mathcal{H}_{B}$). In this case, the control law (8) simplifies to $\dot{\bm{\mathbf{x}}}_{r}(t)=\frac{1}{p(\bm{\mathbf{x}}_{r}(t))}\nabla p(\bm{\mathbf{x}}_{r}(t))$. This non-adaptive approach leads to significantly slower convergence, especially from areas with low signal strength, unlike our method which enables effective SS even with weaker signals. This property of the tuning mechanism is beneficial for SS in unknown environments, as it expedites the process by allowing a lower signal detection threshold for transitioning from exploration to source seeking. This effect is clearly illustrated in two scenarios in Fig. 4 where the underlying probability field is modeled by a log-normal distribution and the convergence to the source is faster using our adaptive algorithm. We choose log normal in this example to demonstrate the effectiveness of our algorithm even in asymmetric signals that closely match real SoIs, such as acoustic signals. Furthermore, Fig. 4 also shows that BEASST maintains convergence under bounded Gaussian noise, demonstrating robustness of the behavioral entropic gradients and confirming the theoretical stability guarantees established in Lemma 1.
1Input: Occupancy map $\mathcal{M}_{D}$, measured signal strength $P_{dBm}(\bm{\mathbf{x}})$, its gradient $\nabla P_{dBm}(\bm{\mathbf{x}})$, number of sources $N_{s}$;
@@ 2 Output: Robot waypoints $\{\bm{\mathbf{x}}_{t+1}\}$;
3 Initialize: $N_{\text{detected}}=0$, $\bm{\mathbf{x}}_{0}$, $\alpha,\beta=1$$\tau$,$t=0$;
4while $N_{\text{detected}}<N_{s}$ do
@@ 3 Initialize: $N_{\text{detected}}=0$, $\bm{\mathbf{x}}_{0}$, $\alpha,\beta=1$$\tau$,$t=0$;
  $p_{t}(\bm{\mathbf{x}})\leftarrow\phi(P_{dBm}(\bm{\mathbf{x}}_{t}))$; // Normalize using Eq. 2
@@ 11    $\bm{\mathbf{x}}_{t+1}\leftarrow\bm{\mathbf{x}}_{t}+\gamma\nabla\log(w(p_{t}(\bm{\mathbf{x}})))$;
13        $N_{\text{detected}}\leftarrow N_{\text{detected}}+1$;
14
15       $(\mathcal{S}_{\text{pending}},\mathcal{S}_{\text{detected}})\leftarrow\text{UpdateSourceSets}(\mathcal{S}_{\text{pending}},\mathcal{S}_{\text{detected}},\bm{\mathbf{x}}_{t})$ ;
@@ 21 $\bm{\mathbf{x}}_{t+1}\leftarrow\text{GPBasedExp}(\mathcal{M}_{D},\text{p(occ(x))})$;
  $\mathcal{M}_{D}$ $\leftarrow\text{Update}$ ($\mathcal{M}_{D}$,p(occ(x))); // SLAM Update [24]
22
@@ V Behavioral Source Seeking with Exploration
The robot initially explores the unknown environment. Upon detecting a signal, above a predefined normalized threshold $\tau>0$, it transitions to SS, driven by the gradients (9) from normalized signal strength $p_{t}(\bm{\mathbf{x}})$. This operation is performed by SourceSensing(.) function in Algorithm 1. The gradient is computed using Gradient(.), which naturally guides the robot towards the strongest perceived signal using adaptive $\alpha_{t}$. For practical implementation, a discrete-time update for the robot’s position is given in Line 10 of Algorithm 1, where $\bm{\mathbf{x}}(t)$ is the desired waypoint for the robot to track. When multiple SoIs are detected, the framework explicitly maintains a set of detected sources, dividing them into two categories: pending ($S_{\textit{pending}}$) and detected ($S_{\textit{detected}}$). Detected but unverified SoIs are stored as ‘virtual frontiers’ (pending sources) for later inspection. Once the robot verifies the currently pursued source, it marks it as detected and removes it from the active list; the corresponding region is masked in the signal map to prevent re-targeting using the UpdateSourceSets. Meanwhile, locations of other significant but unverified signals remain stored as virtual frontiers, serving as waypoints for future exploration in areas with pending sources. This process ensures comprehensive coverage of all detected regions. If no significant signal is detected, the robot prioritizes exploration, moving towards regions of high entropy based on GP-based exploration and Omnimapper [24] updates the map (Update Function). This adaptive switching allows the robot to efficiently locate multiple sources.